How many moles of S are in each of the following quantities?

120 g SO2
1
Expert's answer
2020-10-07T07:32:27-0400


Molar mass of SO2 =( atomic mass of Sulphur + 2 X atomic mass of oxygen)

=( 32 + 2 x 16) g/mol

= 64 g/mol


120 grams of SO2 = ( 120 g /64 g.mol )

= (120/64) mol SO2


1 mole of SO2 contain 1 mole of 'S'

(120/64) mole of SO2 contain = (120/64) mole 'S'

= 1.875 mole 'S'


Hence, 1.875 mole 'S' present in 120 grams of SO2.